the perimeter of a rectangle is 84. The length is 6m more than twice the width. Find the dimensions.
Answer by amarjeeth123(569)

Let the width be x.
Then the length is (2x+6).
Perimeter=2(length+width)
2(2x+6+x)=84
3x+6=42
x+2=14
x=12.
The width of the rectangle is 12.
The length of the rectangle is 30.